# Line-of-sight shear in SLACS strong lenses

### v1

@@ Abstract
Context. Inhomogeneities along the line of sight in strong gravitational lensing distort the images produced, in an effect called shear. If measurable, this shear may provide independent constraints on cosmological parameters, complementary to traditional cosmic shear.
Aims. We model 50 strong gravitational lenses from the Sloan Lens ACS (SLACS) catalogue with the aim of measuring the line-of-sight (LOS) shear for the first time. We use the ‘minimal model’ for the LOS shear, which has been shown to be theoretically safe from degeneracies with lens model parameters, a finding which has been confirmed using mock data.
Methods. We use the dolphin automated modelling pipeline, which uses the lenstronomy software as a modelling engine, to model our selected lenses. We model the main deflector with an elliptical power law profile, the lens light with elliptical Sérsic profiles and the source with a basis set of shapelets and an elliptical Sérsic profile.
Results. We successfully obtain a line-of-sight shear measurement from 18 of the 50 lenses. We find that these LOS shear measurements are consistent with external shears measured in recent works using a simpler shear model, which are larger than those expected from weak lensing. Neglecting the post-Born correction to the potential of the main deflector due to foreground shear leads to a propagation of degeneracies to the LOS shear measurement, and the same effect is seen if a prior is used to connect the lens mass and light ellipticities. The inclusion of an octupole moment in the lens mass profile does not lead to shear measurements that are in better agreement with the expectations from weak lensing.
Key Words.:
Gravitational lensing: strong
@@ 1 Introduction
Gravitational lensing provides a unique window into the cosmology of our Universe on a wide range of scales. The strong lensing regime, in which a single massive deflector lenses a single source, has been used to measure the expansion rate of the Universe at $z=0$, $H_{0}$ (Refsdal, 1964; Wong et al., 2020; Birrer et al., 2020, 2024) via observations of the difference in arrival times of images of variable sources. Strong lensing has been used to probe properties of the lens galaxies, such as their density profiles (Treu et al., 2006; Shajib et al., 2021), and, by studying lens galaxies at different redshifts, their evolution over time (Chae, 2010; Bezanson et al., 2011; Sahu et al., 2024). It has also been used to search for dark matter sub-haloes both within the main deflector halo and along the line of sight between the observer and lens (Vegetti & Koopmans, 2009; Vegetti et al., 2012, 2014; Hezaveh et al., 2016; Nightingale et al., 2024).
@@
The weak lensing regime probes larger scales. Unlike strong lensing, multiple images of the same source are not produced; rather, small distortions are induced on the single image of each source, artificially squashing and aligning them. The leading order shape distortion is shear, and its effect cannot be disentangled from the intrinsic shape of individual galaxies or their intrinsic alignment. The cosmic shear signal is detected by correlating the shapes of millions of galaxies in a given galaxy survey. Once extracted, cosmic shear can be used to place constraints on cosmological parameters like the matter density $\Omega_{\rm m}$ and the clustering parameter $\sigma_{8}$. For a recent review of weak gravitational lensing, we refer the reader to Prat & Bacon (2025).
@@
By virtue of their existence in our inhomogeneous Universe, strong lensing images will also experience weak lensing distortions, with Einstein rings in particular supplying the notion of a ‘standardisable shape’ from which shear may be robustly measured [^1] In practice, shear can be recovered even when the image is not a complete Einstein ring, though the precision of the measurement may be degraded (Hogg et al., 2023)., providing an alternative and independent probe of cosmology to the standard weak lensing shear. This idea was first investigated by Birrer et al. (2017), who modelled the lens COSMOS 0038+4133 with a shear model that consisted of four parameters: two components to describe the foreground shear as projected onto the lens plane, and two components to describe the background shear as projected onto the source plane. Whilst the shear parameters were shown to be constrained with high precision, this analysis revealed strong degeneracies between the shear parameters and the components of the ellipticity of the lens mass. This implies that, with this model, the constraints on the shear cannot be disentangled from the constraints on the lens mass, and the measurements cannot be safely used to obtain cosmological information.
@@
Fleury et al. (2021) developed the minimal line-of-sight shear model to address this problem. Starting with three separate shears, the model exploits a source-position transformation to build a shear parameter, the ‘line-of-sight (LOS) shear’, which at leading order (i.e. for small values of the shear) is mathematically free from degeneracy with any parameter used to describe the lens mass. The model fully describes the problem with the fewest possible parameters, and is therefore called the minimal LOS shear model – we will review the formalism behind this model in more detail below. Thanks to this freedom from degeneracies, it is the LOS shear in the minimal model which should be targeted for eventual cosmological inference. It was shown by Hogg et al. (2023) that the LOS shear is indeed free from degeneracies with lens model parameters using mock strong lensing images. It is therefore timely to attempt to measure this quantity in real strong lens imaging data.
@@
This paper is organised as follows: in section 2 we review strong lensing theory and the minimal model; in section 3 we discuss the data used in this work and the method used to analyse it; in section 4 we present and discuss our results, and we conclude with section 5.
@@ 2 Strong lensing and the line-of-sight shear
In this section we review the strong lensing formalism and the derivation of the minimal model for the LOS shear. For full details, the reader is referred to Fleury et al. (2021).
@@
To a good approximation, further inhomogeneities beyond the main deflector can be treated as tidal perturbations, which induce a shear and convergence on the strong lens image. In this situation, Equation 1 becomes
@@
where the amplification matrices $\mathcal{\bm{A}}_{\rm os}$, $\mathcal{\bm{A}}_{\rm ds}$ and $\mathcal{\bm{A}}_{\rm od}$ encapsulate the LOS perturbations between observer and source, main deflector and source, and observer and main deflector respectively. They can be parameterised as
$\bm{\mathcal{A}}_{\mathrm{ab}}=\bm{1}-\bm{\Gamma}_{\mathrm{ab}}\ ,\quad\bm{\Gamma}_{\mathrm{ab}}=\begin{bmatrix}\kappa_{\mathrm{ab}}+\rm{Re}\left(\gamma_{\mathrm{ab}}\right)&\rm{Im}\left(\gamma_{\mathrm{ab}}\right)-\omega_{\mathrm{ab}}\\ \rm{Im}\left(\gamma_{\mathrm{ab}}\right)+\omega_{\mathrm{ab}}&\kappa_{\mathrm{ab}}-\rm{Re}\left(\gamma_{\mathrm{ab}}\right)\end{bmatrix},$ (6)
where $\mathrm{ab}\in\{\mathrm{os},\mathrm{ds},\mathrm{od}\}$ and $\kappa_{\rm ab}$ is the convergence, $\gamma_{\rm ab}$ the shear and $\omega_{\rm ab}$ the rotation of an image as seen at position a with respect to its source at b, produced by lens-lens coupling between the main deflector and the perturbers. A point regarding notation: we later denote the real component of any shear with $\gamma_{1}$, and the imaginary component with $\gamma_{2}$, with the magnitude $\gamma=\sqrt{\gamma_{1}^{2}+\gamma_{2}^{2}}$.
Measuring the three individual shears $\gamma_{\rm os}$, $\gamma_{\rm ds}$ and $\gamma_{\rm od}$ from a strong lensing image is impossible, since they are degenerate with each other and with parameters of the model used to describe the gravitational potential of the main deflector, $\psi(\bm{\theta})$ (Hogg et al., 2023); if the iso-potential contours are elliptical, there is an exact degeneracy between the ellipticity and the foreground shear $\gamma_{\rm od}$. This is due to the fact that we have no access to the true, unlensed source position $\bm{\beta}$ or its unlensed morphology, meaning that up to linear transformations of the source, image positions are preserved; this is called the source-position transformation degeneracy (Schneider & Sluse, 2014), which is a generalisation of the mass-sheet degeneracy (Falco et al., 1985).
However, this degeneracy can also be exploited in order to define a shear quantity which is measurable and free from degeneracies with lens model properties. This can be done by arbitrarily fixing a definition for $\bm{\beta}$, analogous to lifting a gauge freedom by fixing the gauge. It was shown by Fleury et al. (2021) that a minimal model for LOS perturbations can be found by multiplying Equation 5 by $\bm{\mathcal{A}}_{\mathrm{od}}\bm{\mathcal{A}}^{-1}_{\mathrm{ds}}$, yielding
@@
It was argued by Fleury et al. (2021) that the LOS shear, the combination $\gamma_{\rm LOS}\equiv\gamma_{\rm od}+\gamma_{\rm os}-\gamma_{\rm ds}$ which arises from the $\bm{\mathcal{A}}_{\mathrm{LOS}}$ amplification matrix in the minimal model, is not degenerate with the ellipticity of the main lens at leading order. This was further investigated by Hogg et al. (2023), who fit different models to a Hubble Space Telescope-like catalogue of mock strong lensing images, showing that $\gamma_{\rm LOS}$ is recoverable and non-degenerate with other lens model parameters, provided the lens mass model has a complexity equivalent to that which the mock images were generated with. The current work takes the logical next step, providing the first measurement of the LOS shear in real observational data, laying the groundwork for the cosmological constraints described above.
@@ 3.1 The strong lens sample
The Sloan Lens ACS (SLACS) strong lens catalogue is a large, well-studied sample of lenses, making it an ideal testbed for a first measurement of the LOS shear. Candidate lenses were identified spectroscopically in Sloan Digital Sky Survey (SDSS) data and followed up with Hubble Space Telescope (HST) observations (Bolton et al., 2006). The lenses we model in this work were imaged either with the Advanced Camera for Surveys (ACS) in the F [^5] W filter or with the Wide Field and Planetary Camera 2 (WFPC2) in the F606W filter. The imaging data were obtained under the HST General Observers programmes 10494 (PI: Koopmans), 10798 (PI: Bolton), 10886 (PI: Bolton), and 11202 (PI: Koopmans).
A subset of 50 representative strong lenses from the SLACS catalogue were selected for study by Shajib et al. (2021), with the criteria that the chosen lenses should not have satellite galaxies or complex source morphologies, in order to allow an automated modelling procedure to be applied. Furthermore, lenses were selected that have imaging data in the F [^5] W and F606W filters, as the deflector light tends to be fainter in these bands compared to the lensed source light, which makes it easier to deblend the source and deflector lights during lens modelling. Lastly, no disc-like lenses were selected.
The WFPC2 images were reduced for the original SLACS analysis; ACS images were reduced by Shajib et al. (2021) using the Astrodrizzle package (Avila et al., 2015), and the point-spread function (PSF) for each filter and camera combination were computed using TinyTim (Krist et al., 2011). In this work, we make use of these imaging data and PSFs, modelling the same 50 lenses as Shajib et al. (2021), thus enabling a direct comparison of our results with that previous work.
@@ 3.2 Mass and light models
We model the mass of the main deflector using the elliptical power law (EPL) profile of Tessore & Metcalf (2015). The convergence of this profile is given by
@@
where $\theta_{\rm E}$ is the Einstein radius, $\gamma^{\rm EPL}$ is the slope of the power law describing the three-dimensional mass distribution and $q$ is the axis ratio of the ellipse. Whilst the EPL profile is a relatively simplistic choice to describe the lens mass, we use it in order to facilitate comparison of our results with other works that have also used it.
We model the light of the main deflector with two elliptical Sérsic profiles (Sérsic, 1963, 1968). The surface brightness of a Sérsic profile is given by
@@
We use two Sérsic profiles rather than one to model the main deflector’s light, as it has been shown, when fitting the specific lens RXJ1131-1231, that a single Sérsic profile results in significant residual differences between observation and model in the centre of the image (Claeskens et al., 2006; Suyu et al., 2013). We constrain the profiles so that their centres share the same position in the lens plane, and furthermore, fix the Sérsic indices of the two profiles to one and four, $n_{\rm S}=1$ and $n_{\rm S}=4$, representing an exponential and a de Vaucouleurs profile respectively (de Vaucouleurs, 1948).
We model the light of the source with a further two light profiles: an elliptical Sérsic profile plus a set of shapelets. Shapelets are two-dimensional Gauss–Hermite polynomials, and thanks to their forming a complete basis set, any image can be decomposed into a linear combination of these polynomials (Refregier, 2003). They have been used to measure weak lensing shear in galaxy images by Refregier & Bacon (2003), and for numerous source reconstructions from strong lensing images, beginning with Birrer et al. (2015).
We here summarise the shapelets formalism as presented in Refregier (2003). The dimensionless basis functions, or shapelets, are defined in one dimension as
@@
In practice, the parameters of our set of shapelets are the maximum order for the Hermite polynomial $n_{\rm max}\geq n_{1}+n_{2}$, which in turn determines the number of shapelets in the set, $N_{\rm shapelets}=(n_{\rm max}+1)(n_{\rm max}+2)/2$, and the characteristic scale $\beta$. In our analysis, we fix $n_{\rm max}=6$, whilst $\beta$ is inferred by the Markov chain Monte Carlo (MCMC) procedure described below. The centre of the set of shapelets is constrained to be at the same position in the source plane as the centre of the elliptical Sérsic profile which makes up the rest of the source light model.

### v2

@@ Abstract
Inhomogeneities along the line of sight in strong gravitational lensing distort the images produced, in an effect called shear. If measurable, this shear may provide independent constraints on cosmological parameters, complementary to traditional cosmic shear. We model 23 strong gravitational lenses from the Sloan Lens ACS (SLACS) catalogue with the aim of measuring the line-of-sight (LOS) shear for the first time. We use the ‘minimal model’ for the LOS shear, which has been shown to be theoretically safe from degeneracies with lens model parameters, a finding which has been confirmed using mock data. We use the dolphin automated modelling pipeline, which uses the lenstronomy software as a modelling engine, to model our selected lenses. Across the 23 lenses, we measure the LOS shear with a mean magnitude of $0.056\pm 0.013$. Neglecting the post-Born correction to the potential of the main deflector due to foreground shear leads to a propagation of degeneracies in the LOS shear measurement with other lens model parameters, and the inclusion of an octupole moment in the lens mass profile does not lead to shear measurements that are in better agreement with the expectations from weak lensing.
keywords:
Strong gravitational lensing, cosmology
thanks: natalie.hogg@ast.cam.ac.uk
@@ 1 Introduction
Gravitational lensing provides a unique window into the cosmology of our Universe on a wide range of scales. The strong lensing regime, in which a single massive deflector lenses a single source, has been used to measure the expansion rate of the Universe at $z=0$, $H_{0}$ (Refsdal1964b; Wong:2019kwg; Birrer:2022chj; TDCOSMO:2025dmr) via observations of the difference in arrival times of images of variable sources. Strong lensing has been used to probe properties of the lens galaxies, such as their density profiles (Treu2006; Shajib2021; Shajib2024), and, by studying lens galaxies at different redshifts, their evolution over time (Chae2010; Bezanson2011; Sahu:2024ayh; Sheu:2024sjd). It has also been used to search for dark matter sub-haloes both within the main deflector halo and along the line of sight between the observer and lens (Vegetti_2009; Vegetti_2012; Vegetti_2014; Hezaveh:2016ltk; Nightingale2024b).
@@
The weak lensing regime probes larger scales. Unlike strong lensing, multiple images of the same source are not produced; rather, small distortions are induced on the single image of each source, artificially squashing and aligning them. The leading order shape distortion is shear, and its effect cannot be disentangled from the intrinsic shape of individual galaxies or their intrinsic alignment. The cosmic shear signal is detected by correlating the shapes of millions of galaxies in a given galaxy survey. Once extracted, cosmic shear can be used to place constraints on cosmological parameters like the matter density $\Omega_{\rm m}$ and the clustering parameter $\sigma_{8}$. For a recent review of weak gravitational lensing, we refer the reader to Prat:2025ucy.
@@
By virtue of their existence in our inhomogeneous Universe, strong lensing images will also experience weak lensing distortions, with Einstein rings in particular supplying the notion of a ‘standardisable shape’ from which shear may be robustly measured [^1] In practice, shear can be recovered even when the image is not a complete Einstein ring, though the precision of the measurement may be degraded (Hogg:2022ycw)., providing an alternative and independent probe of cosmology to the standard weak lensing shear.
There is a twofold benefit to using strong lensing shear data in cosmological analyses: firstly, simply by virtue of enlarging the data vector, there will be an automatic increase in statistical power when constraining cosmological parameters with lensing data; and secondly, whilst systematic uncertainties will undoubtedly be present in such constraints, they will nonetheless be complementary, rather than additive, to the systematics present in standard weak lensing data, such as galaxy bias and intrinsic alignments. The most obvious systematic, shear and lens mass model choice and their subsequent effect on shear measurements, are the focus of this work.
The idea of using strong lensing images to measure shear for cosmology was first investigated by Birrer:2016xku, who modelled the lens COSMOS 0038+4133 with a shear model that consisted of four parameters: two components to describe the foreground shear as projected onto the lens plane, and two components to describe the background shear as projected onto the source plane. Whilst the shear parameters were shown to be constrained with high precision, this analysis revealed strong degeneracies between the shear parameters and the components of the ellipticity of the lens mass. This implies that, with this model, the constraints on the shear cannot be disentangled from the constraints on the lens mass, and the measurements cannot be safely used to obtain cosmological information.
@@
Fleury:2021tke developed the minimal line-of-sight shear model to address this problem. Starting with three separate shears, the model exploits a source-position transformation to build a shear parameter, the ‘line-of-sight (LOS) shear’, which is mathematically free from degeneracy with ellipticity in the lens mass. The model fully describes the problem with the fewest possible parameters, and is therefore called the minimal LOS shear model – we will review the formalism behind this model in more detail below. Thanks to this freedom from degeneracies, it is the LOS shear in the minimal model that should be targeted for eventual cosmological inference. It was shown by Hogg:2022ycw that the LOS shear is indeed free from degeneracies with lens model parameters using mock strong lensing images. It is therefore timely to attempt to measure this quantity in real strong lens imaging data.
A further motivation of this work is to explore the effect of different shear and mass model parametrisations, in light of the findings of Etherington:2023yyh, who showed that external shear magnitudes in some strong lenses are suspiciously large. They proposed that unmodelled complexity in the lens mass may be the reason for this. We therefore explore the role of foreground shear in the minimal LOS model, as well as octupolar distortions in the lens mass, in the sample of lenses studied in this work.
@@
This paper is organised as follows: in section 2 we review strong lensing theory and the minimal model; in section 3 we discuss the data used in this work and the method used to analyse it; in section 4 we present and discuss our results, namely fitting our sample of strong lenses with two different shear parametrisations and two different mass model parametrisations, including an octupolar distortion; we conclude with section 5.
@@ 2 Strong lensing and the line-of-sight shear
In this section we review the strong lensing formalism and the derivation of the minimal model for the LOS shear. For full details, the reader is referred to Fleury:2021tke.
@@
To good approximation, further inhomogeneities beyond the main deflector can be treated as tidal perturbations, which induce a shear and convergence on the strong lens image. The exceptions to this are, for example, double plane lensing (Gavazzi:2008aq; Collett:2020lii; Dux:2024vvq) or systems for which nearby satellite galaxies are explicitly included in the model, e.g. Nightingale:2022bhh. Non-tidal perturbations will also induce flexion on strong lensing images (Goldberg:2004hh; Bacon:2005qr; Duboscq:2024asf).
In the case of single plane lensing with tidal perturbations, Equation 1 becomes
@@
where the amplification matrices $\bm{\mathcal{A}}_{\mathrm{os}}$, $\bm{\mathcal{A}}_{\mathrm{ds}}$ and $\bm{\mathcal{A}}_{\mathrm{od}}$ encapsulate the LOS perturbations between observer and source, main deflector and source, and observer and main deflector respectively. They can be parameterised as
$\bm{\mathcal{A}}_{\mathrm{ab}}$ $=\bm{1}-\bm{\Gamma}_{\mathrm{ab}}\ ,$ (6)
$\bm{\Gamma}_{\mathrm{ab}}$ $=\begin{bmatrix}\kappa_{\mathrm{ab}}+\rm{Re}\left(\gamma_{\mathrm{ab}}\right)&\rm{Im}\left(\gamma_{\mathrm{ab}}\right)-\omega_{\mathrm{ab}}\\ \rm{Im}\left(\gamma_{\mathrm{ab}}\right)+\omega_{\mathrm{ab}}&\kappa_{\mathrm{ab}}-\rm{Re}\left(\gamma_{\mathrm{ab}}\right)\end{bmatrix}$, (7)
where $\mathrm{ab}\in\{\mathrm{os},\mathrm{ds},\mathrm{od}\}$ and $\kappa_{\rm ab}$ is the convergence, $\gamma_{\rm ab}$ the shear and $\omega_{\rm ab}$ the rotation of an image as seen at position ‘a’ with respect to its source at ‘b’, produced by lens–lens coupling between the main deflector and the perturbers. A point regarding notation: we later denote the real component of any shear with $\gamma_{1}$, and the imaginary component with $\gamma_{2}$, with the magnitude $|\gamma|=\sqrt{\gamma_{1}^{2}+\gamma_{2}^{2}}$.
Measuring the three individual shears $\gamma_{\rm os}$, $\gamma_{\rm ds}$ and $\gamma_{\rm od}$ from a strong lensing image is impossible, since they are degenerate with each other and with parameters of the model used to describe the gravitational potential of the main deflector, $\psi(\bm{\theta})$ (Hogg:2022ycw); if the iso-potential contours of the lens mass are elliptical, there is an exact degeneracy between the ellipticity and the foreground shear $\gamma_{\rm od}$. This is due to the fact that we have no access to the true, unlensed source position $\bm{\beta}$ or its unlensed morphology, meaning that up to linear transformations of the source, image positions are preserved; this is called the source-position transformation degeneracy (Schneider:2013wga), which is a generalisation of the mass-sheet degeneracy (Falco1985).
However, this degeneracy can also be exploited in order to define a shear quantity which is measurable and free from degeneracies with lens model properties. This can be done by arbitrarily fixing a definition for $\bm{\beta}$, analogous to lifting a gauge freedom by fixing the gauge. It was shown by Fleury:2021tke that a minimal model for LOS perturbations can be found by multiplying Equation 5 by $\bm{\mathcal{A}}_{\mathrm{od}}\bm{\mathcal{A}}^{-1}_{\mathrm{ds}}$, yielding
@@
It was argued by Fleury:2021tke that the LOS shear, the combination $\gamma_{\rm LOS}\equiv\gamma_{\rm od}+\gamma_{\rm os}-\gamma_{\rm ds}$ which arises from the $\bm{\mathcal{A}}_{\mathrm{LOS}}$ amplification matrix in the minimal model, is not degenerate with the ellipticity of the main lens at leading order. This was further investigated by Hogg:2022ycw, who fit different models to a Hubble Space Telescope-like catalogue of mock strong lensing images, showing that $\gamma_{\rm LOS}$ is recoverable and non-degenerate with other lens model parameters, provided the lens mass model has a complexity equivalent to that with which the mock images were generated. The current work takes the logical next step, providing the first measurement of the LOS shear in real observational data and beginning to explore potential systematics in that measurement due to the shear and lens mass models, laying the groundwork for the cosmological constraints described above.
@@ 3.1 The strong lens sample
The Sloan Lens ACS (SLACS) strong lens catalogue is a large, well-studied sample of lenses, making it an ideal testbed for a first measurement of the LOS shear. Candidate lenses were identified spectroscopically in Sloan Digital Sky Survey (SDSS) data and followed up with Hubble Space Telescope (HST) observations (Bolton2006). The lenses we model in this work were imaged either with the Advanced Camera for Surveys (ACS) in the F [^5] W filter or with the Wide Field and Planetary Camera 2 (WFPC2) in the F606W filter. The imaging data were obtained under the HST General Observers programmes 10494 (PI: Koopmans), 10798 (PI: Bolton), 10886 (PI: Bolton), and 11202 (PI: Koopmans).
A subset of 50 representative strong lenses from the SLACS catalogue were selected for study by Shajib2021, with the criteria that the chosen lenses should not have satellite galaxies or complex source morphologies, in order to allow an automated modelling procedure to be applied. Furthermore, lenses were selected that have imaging data in the F [^5] W and F606W filters, as the deflector light tends to be fainter in these bands compared to the lensed source light, which makes it easier to deblend the source and deflector lights during lens modelling. Lastly, no disc-like lenses were selected. In their study, 23 of the 50 were modelled successfully.
The WFPC2 images were reduced for the original SLACS analysis; ACS images were reduced by Shajib2021 using the Astrodrizzle package (Avila2015), and the point-spread function (PSF) for each filter and camera combination were computed using TinyTim (Krist2011). In this work, we make use of these imaging data and PSFs, modelling the same 23 lenses as Shajib2021. These lenses are listed in Table 1. We further extend our sample to 45 lenses in Hogg2025c.
@@ 3.2 Mass and light models
We model the mass of the main deflector using the elliptical power law (EPL) profile of Tessore:2015baa. The convergence of this profile is given by
@@
where $\theta_{\rm E}$ is the Einstein radius, $\gamma^{\rm EPL}$ is the slope of the power law describing the three-dimensional mass distribution and $q$ is the axis ratio of the ellipse. The EPL profile is a relatively simplistic description of the lens mass, and suffers from known systematics in the absence of additional complexity (Schneider2013; Cao2022). Nonetheless, it has been widely used in relevant literature, e.g. Etherington:2022nzt; Etherington:2023yyh, and thus serves as a foundational model for comparisons with such works, and one to which additional flexibility can be added.
The Einstein radius of each lens is not included in our Markov chain Monte Carlo (MCMC) parameter inference – which we describe in more detail in subsection 3.4 – as a sampled parameter. Instead, we fix $\theta_{\rm E}$ of each lens to its best fit value as reported by Shajib2021 (see Table 1 of that work). The motivation for this choice is to improve the computational efficiency of the MCMC. We note that this choice may lead to some loss of uncertainty in the remaining parameters, but since $\theta_{\rm E}$ typically has the smallest inherent uncertainty of all parameters in the model, the change is not significant.
We model the light of the main deflector with two elliptical Sérsic profiles (Sersic1963; Sersic1968). The surface brightness of a Sérsic profile is given by
@@
We use two Sérsic profiles rather than one to model the main deflector’s light, as it has been shown, when fitting the specific lens RXJ1131-1231, that a single Sérsic profile results in significant residual differences between observation and model in the centre of the image (Claeskens:2006ky; Suyu2013). We constrain the profiles so that their centres share the same position in the lens plane, and furthermore, fix the Sérsic indices of the two profiles to one and four, $n_{\rm S}=1$ and $n_{\rm S}=4$, representing an exponential and a de Vaucouleurs profile respectively (deVaucouleurs1948).
We model the light of the source with a further two light profiles: an elliptical Sérsic profile plus a set of shapelets. Shapelets are two-dimensional Gauss–Hermite polynomials, and thanks to their forming a complete basis set, any image can be decomposed into a linear combination of these polynomials (Refregier:2001fd). They have been used to measure weak lensing shear in galaxy images by Refregier:2001fe, and for numerous source reconstructions from strong lensing images, beginning with Birrer:2015rpa.
We here summarise the shapelets formalism as presented in Refregier:2001fd. The dimensionless basis functions, or shapelets, are defined in one dimension as
@@
In practice, the parameters of our set of shapelets are the maximum order for the Hermite polynomial $n_{\rm max}\geq n_{1}+n_{2}$, which in turn determines the number of shapelets in the set, $N_{\rm shapelets}=(n_{\rm max}+1)(n_{\rm max}+2)/2$, and the characteristic scale $\beta$. In our analysis, we fix $n_{\rm max}$ to a specific value for each lens, starting from the values used in Shajib2021 and updating some in a trial-and-error process. The values used are listed in Table 1. The scale parameter $\beta$ is inferred by the MCMC procedure described below. The centre of the set of shapelets is constrained to be at the same position in the source plane as the centre of the elliptical Sérsic profile that makes up the rest of the source light model.
Gaussian shapelets suffer from the limitation that the 0th order function – the Gaussian – is not a realistic representation of the light profile of a typical galaxy (Ngan:2008rg). This can lead to over-fitting, particularly in the outskirts of the galaxy. As an alternative, Berge:2019nyt proposed one and two-dimensional exponential shapelets, based on wave functions of the hydrogen atom, which they argued are more suitable for describing galaxy shapes in astronomical imaging, amongst other applications. In this work, we tested the exponential shapelets for source reconstruction on five of our sample of lenses, finding generally poorer fits. For this reason, and for the reason of preserving a fair comparison between our results and that of Shajib2021, we continued to use the Gaussian shapelets for our source reconstructions. A more rigorous exploration of the use of exponential, or other forms of shapelets, is left to a future work.